### v1

@@ 6.5 Example Application Scenario for the Reference Architecture
    A. Modelling notation for quantum services: As per Figure 8, we have used the UML as the modelling notation to create the service model based on UML component diagram and UML sequence diagram. To create quantum-enabled models for QSE, the quantum UML profile extends the traditional UML diagrams to support structural and behavioral modeling of quantum software [52]. Specifically, the UML component diagram in Figure 8 shows a structural view of the quantum services and their interconnection. Following the notations from UML profile [63], from a service modelling perspective each service is represented as an individual component. The component provides a computational service (e.g., generate a random number) and communicates with another component(s) using a connector. For example, the service named GetGCD interconnects with another service named Controller to generate the Greatest Common Divisor (GCD) of a randomly generated number as part of the algorithmic implementation. In contrast to a structural view of the quantum service model using a component diagram, the UML sequence diagram reflects the behavior of the services in terms of message passing among the services to enable service communication. For example, in the sequence diagram, the controller service C: Controller passes a message Generate(N) to the number generator service N: NumGenerator that generates a random number R and returns it to the controller service. A number of other UML diagrams can also be used, such as the UML use case diagram to represent the QSRs. For demonstration purposes, we have used the UML component and sequence diagrams to exemplify service modelling in terms of structural composition and behavioural representation of the services to be assembled.
@@
    B. Pattern-based modelling: Service design and development patterns provide reuse knowledge and best practices to engineer service-oriented solutions [26]. In Figure 8, we have applied two patterns namely the orchestrator and classic-quantum split pattern. The orchestrator pattern is a classic SOA pattern that helps to orchestrate the execution of a number of services to complete a service-driven task. For example, the application of the orchestrator pattern helped to orchestrate the factorisation of a randomly generated prime number via NumGenerator and Factorise services. The classic-quantum split pattern helps to split the functionality between quantum and classic computing, also referred to as hybrid quantum computing. In the context of Figure 8, the Controller service orchestrates the functionality between classical and quantum microservices. For example, the services named NumGenerator and GetGCD can be executed on a classical machine, whereas the services QunatumModularExponentiation, QunatumInverseQFT, and Factorise are executed on the quantum machine. As part of the reference architecture, the modeling phase exploits modeling notation and applies a pattern to create a service design, i.e., a ’visual model’ that can be assembled into executable quantum microservices.
@@
  • Phase III - Service Assembly: Service assembly refers to assembling a service, i.e., identifying a use-case and programming the service that implements the usecase as shown in Figure 8 and detailed below.
@@
    A. Service usecase: Implementation details are elaborated based on Figure 8 that demonstrates the conception, modelling, assembly, and deployment of software services that implement the Shor’s. Shor’s algorithm is a quantum computing algorithm that is used for factoring integers in polynomial time [64]. Shor’s algorithm has a number of use cases such as prime factorisation, cryptography (e.g., breaking the RSA scheme), and finding the period of a function. In the scope of this work, we only focus on developing microservices, executed on a QC platform that implement Shor’s algorithm for prime factorisation. We provide an illustrative case in Figure 8 that exemplifies the architecture-based development of the services (from Figure 7) to implement Shor’s algorithm. Specifically, Figure 7 as a reference architecture highlights what needs to be done? by providing a blue-print to architect quantum service-orientation.
@@
    B. Service programming: In order to implement the use-case service needs to be programmed for its execution. This means that design specifications from the modeling phase need to be translated into executable specifications as part of service assembly. Figure 8, shows the algorithmic design and its implementation as part of service assembly. The algorithm shows a partial assembly of services by exemplifying the classical services that generate a random number and compute the GCD for quantum computation of the Shor’s algorithm. The UML sequence diagram is used as a model to assemble the service via an algorithm and its implementation in a given programming language. A synoptic view of the code snippet (written in C#) is shown that highlights the source code skeleton for a controller service PrimeFactorisationController that orchestrates the classical and quantum services. Figure 8 implements the reference architecture to highlight how it is to be done? with an illustrative example of implementing quantum service-orientation.
@@
  • Phase IV - Service Deployment as the last phase in the reference architecture, deployment consists of two main activities namely service hosting and service execution. While execution and hosting primarily depend on the QC platform on which the quantum services are deployed.
@@
    Service Hosting Platform: The hosting of the assembled service is presented as a UML deployment diagram as in Figure 8. UML deployment diagram show two deployment nodes (hosting machines). The classical compute node hosts three services NumGenerator, GetGCD, and Controller, whereas the quantum compute node hosts three services, namely QunatumModularExponentiation, QunatumInverseQFT, and Factor.
    Figure 8 b) shows the user interface of QADL, a prototype that we developed to model and assemble, i.e., represent and formally describe quantum software components that can provide computational services. In Figure 8 b), left-side, the QADL script specifies the architectural description, declaring QuBits and applying the QUGates (Hadamard, Pauli-X, and CNOT), followed by measurement. This description corresponds to the core quantum subroutine of Shor’s algorithm - superposition (via Hadamard), controlled operations for modular exponentiation (via CNOT), and state inversion (via Pauli-X) - leading to the measurement for computing the periodicity. QADL script for architectural specification of Shor’s algorithm via a quantum circuit diagram is presented in Appendix A - Listing 2.
@@
    In Figure 8 b), towards the right, QADL represents qunatum circuit diagram to model the algorithm. The description and representation with QADL enables Shor’s algorithm to be modeled, visualised, and executed as a service abstraction that supports design at higher-level of software abstraction rather than low-level implementations. Low-level details, such as circuit creation and component execution within QADL is enabled via IBM Qiskit - an open-source QSW development framework created by IBM [65]. The Qiskit provides the tools to design, simulate, and run quantum programs on both simulators and real quantum hardware, its primary role is to provide a modular and descriptive framework for defining QSW.
7 Emerging Trends of Research on QCaaS (RQ-3)
@@
We now answer RQ-3 that aims to report the emerging trends, as dimensions of potential future research, on engineering software services for quantum computing. To identify these trends, we specifically reviewed the details about the objectives, evaluation/demonstrations, as well as limitations and future research details in each reviewed study. This helped us to identify a number of trends that can be generally classified into five (5) main areas, detailed in this section. The identified trends presented in this section do not represent a comprehensive list, instead, it highlights some possible future research based on the data from existing literature. To maintain a consistent presentation of results, we have visualised the emerging trends identified in Figure 9 according to the service lifecycle activities [12] [26] as in Table 2. For example, Figure 9 highlights that during the conception of quantum services, an empirical identification and thematic classification of Quantum Significant Requirements (QSRs) is an area of future research. QSRs include quality aspects that complement the functional aspects to ensure the required functionality and desired quality of QCaaS as part of quantum domain engineering activity in QSE.
7.1 Process-centric and Human-assisted QCaaS Development
@@
Process-centred engineering is manifested in service life cycle to structure a multitude of activities, such as service design, development, and delivery, into a coherent process that can be executed in an incremental fashion [22]. Although the evidence from the selected studies is accumulated and organised under SOA life cycle, however, as an individual contribution, the studies lacked a process-centric approach and highlighted the needs for processes attuned to the needs of QSE [22]. These processes, such as Quantum DevOps, Quantum microservicing or agile methods for quantum service/software etc., are seen as quantum specific genre of SE processes that align with the needs of QSE [66]. For example, in the agile method for QSE, quantum domain engineering can help accumulate the domain information of a quantum system (hardware/software) to develop a design model that can act as a blueprint to implement quantum software and services. Process-centric approaches can also support tools (for automation) and professional roles (human decision support) to engineer quantum services. We identified the need for human roles as QSE professionals to effectively undertake QCaaS development.
@@
Human roles can enrich the process - synergising knowledge of quantum physics practices of software engineering - to support activities of quantum domain engineering, software architecting, and simulation management as expertise that currently lacks in quantum service-orientation.
7.2 Empirical Identification and Classification of Quantum Significant Requirements
@@
The concept of Quantum Significant Requirements (QSRs) can be traced to well-established role of Architecturally Significant Requirements (ASRs) in software and service engineering [48]. QSRs as part of the quantum domain engineering activity can help elicit and document the functional and non-functional aspects (also quality aspects or attributes) of quantum software. Existing research primarily focuses on functional aspects of services while overlooking the non-functional aspects that include but are not limited to QuBit utilisation, energy efficiency, quantum vendor lock-ins, QPU elasticity, quantum error mitigation, that impact the development and operationalisation or QCaaS solutions. For example, during the quantum domain engineering activity the hardware aspects (operations of QuGates) can be mapped to software aspects (components and connectors) that can help with splitting the computation tasks between classical and quantum computers using the classic-quantum split pattern.
@@
Empirical analysis of available data available via the social coding platforms (e.g., GitHub) grounds theoretical QSRs identified via literary analysis, like Qubit utilization and error mitigation, in a practical and developer-centric context. By examining real-world projects, such as those on GitHub, research can validate or refine the identified requirements with actual developer priorities. This evidence-based approach transforms QSRs from abstract concepts into a practical taxonomy reflective of real quantum software development needs [24]. Furthermore, empirical identification and classification of the QSRs can lead to the creation of a comprehensive QSR taxonomy. By combining the findings of the literature review with data mined from open-source repositories can capture both the existing and emergent, implicit requirements from developer artifacts like issue trackers of qunatum software systems. Classifying these findings allows for the development of a taxonomical classification of QSRs and identifies critical trade-offs, such as between quantum circuit depth and quantum error mitigation [22].
@@
Quantum significant requirements are rooted in the concept of ASRs that aims to identify, classify, and document functional and quality aspects of quantum software services. QSRs for quantum services can complement the functional aspects with quality requirements to ensure the required functionality and desired quality of service. Empirical mining of open-source social doing platforms can validate theoretical QSRs and transform them into a practical, evidence-based taxonomy reflective of real development challenges.
@@ 7.3 Model-driven Quantum Software Servicing
Model-driven Service Engineering (MDSE) enables software engineers and architects to rely on models that can abstract complex and implementation-specific details with human comprehensible visual notations to design and implement software services. Specifically, by exploiting MDSE, software engineers can apply the model transformation to transform the design models into implementation (source code) and validation (test case) models. Modeling notations, such as Service-oriented architecture Modeling Language (SoaML) and more specifically the Q-UML specification [52], provide a metamodel and a UML profile for the specification and design of services within a service-oriented architecture. MDSE can benefit novice developers to to map the flow of algorithms and modules of source code to graphical models for low-code (model-driven) development of quantum services.
Model-driven Quantum Service Development can leverage the principle of MDSE and modelling notations like SoaML and Q-UML to abstract implementation-level complexities with design-level models driven by QSRs. To exploit model-driven development, there is a need for tool support that could enable automation (e.g., model transformation tools) and human decision support (e.g., quantum software architects) to quantum software servicing using MDSE.
@@
7.4 Empiricism in Mining Quantum Service Patterns
@@
Pattern-based software service engineering relies on architectural design and implementation strategies and best practices that can be reused to deliver software services [26]. Existing solutions employ a number of pattern-based solutions, such as classic-quantum split and service wrapping pattern, however, there is no evidence on an empirical discovery of patterns and tactics as reusable knowledge [67]. A lack of empiricism in discovering new patterns hinders the reusability of service design and implementation knowledge during quantum service engineering. One possible dimension for pattern discovery is mining software repositories or social coding platforms (e.g., GitHub) that contain raw knowledge that can be mined as patterns. Pattern-based solutions could complement human expertise with available best practices for service design and implementation. More specifically, novice developers of quantum software services can rely on patterns that abstract complexities of quantum source coding with modularisation of pattern-based solutions.
@@
Pattern discovery via empirically-grounded methods, i.e., mining repositories or social coding platforms of quantum software development, can help leverage reusable design rationale for quantum service engineering. Pattern languages can empower the role of quantum software architects and algorithm designers to rely on reusable knowledge and best practices as opposed to ad-hoc and once-off solutions.
7.5 Continuous Testing and Delivery of Quantum Services

### v2

@@ A. Modelling notation for quantum services:
    As shown in Figure 8, UML is used as the modeling notation to create the quantum service model, employing UML component and sequence diagrams. To support quantum-enabled modeling in QSE, the quantum UML profile extends traditional UML diagrams for both structural and behavioral modeling of quantum software [53]. The UML component diagram in Figure 8 provides a structural view of quantum services and their interconnections. According to the UML profile [66], each service is represented as a separate component that offers a computational function (e.g., generating a random number) and communicates with other components via connectors. For instance, the GetGCD service connects with the Controller service to compute the Greatest Common Divisor (GCD) of a randomly generated number as part of the algorithm. In contrast, the UML sequence diagram captures the behavioral view, illustrating message passing between services to enable communication. For example, in the sequence diagram, the controller service C:Controller sends a Generate(N) message to the number generator service N:NumGenerator, which generates a random number R and returns it to the controller. Other UML diagrams, such as use case diagrams, can also represent QSRs. For demonstration, we used component and sequence diagrams to exemplify service modeling, highlighting both the structural composition and behavioral interactions of services to be assembled.
@@
    B. Pattern-based modelling: Service design and development patterns provide reusable knowledge and best practices for engineering service-oriented solutions [28]. In Figure 8, we applied two patterns: Orchestrator and Quantum-Classic Split. The Orchestrator pattern, a standard SOA pattern, manages the execution of multiple services to complete a service-driven task. For example, it coordinates the factorisation of a randomly generated prime number through the NumGenerator and Factorise services. The Classic-Quantum Split pattern divides functionality between classical and quantum computing, enabling hybrid quantum computing. In Figure 8, the Controller service orchestrates operations between classical and quantum microservices: NumGenerator and GetGCD run on a classical machine, while QuantumModularExponentiation, QuantumInverseQFT, and Factorise execute on the quantum machine. Within the reference architecture, the modeling phase leverages both modeling notations and patterns to create a service design – a visual model that can be assembled into executable quantum microservices.
@@
  • Phase III - Service Assembly: Service assembly involves assembling a service by identifying a use case and developing the corresponding program that implements it, shown in Figure 8 and described below.
@@
    A. Service usecase: The implementation details based on Figure 8, which illustrates the conception, modeling, assembly, and deployment of software services implementing Shor’s algorithm. Shor’s algorithm is a quantum computing algorithm for factoring integers in polynomial time [30], with applications including prime factorisation, cryptography (e.g., breaking RSA), and period finding for functions. In this study, we focus specifically on developing microservices that run on a quantum computing platform to implement Shor’s algorithm for prime factorisation. Figure 8 provides an illustrative example of how services—modeled using the reference architecture in Figure 7 can be developed to realize Shor’s algorithm. The reference architecture in Figure 7 serves as a blueprint, outlining the necessary steps and structure to architect a quantum service-oriented solution.
@@
    B. Service implementation framework: To implement the use case, the service must be programmed for execution, meaning the design specifications from the modeling phase are translated into executable specifications during service assembly. Figure 8 illustrates this process, showing both the algorithmic design and its implementation as part of service assembly. The example demonstrates a partial assembly of services, including classical services that generate a random number and compute the GCD, which support the quantum computation in Shor’s algorithm. A UML sequence diagram serves as the model to guide assembly, mapping the algorithm to its implementation in a programming language. A C# code snippet provides a summarised view of the PrimeFactorisationController service, which orchestrates both classical and quantum services. Figure 8 effectively demonstrates how the reference architecture can be applied in practice, providing a concrete example of implementing quantum service-orientation.
@@
  • Phase IV - Service Deployment As the final phase in the reference architecture, deployment comprises of service hosting and service execution. Both activities largely depend on the quantum computing platform where the quantum services being deployed.
    Service Hosting Platform: The hosting of the assembled services is depicted using a UML deployment diagram in Figure 8. The diagram shows two deployment nodes representing the hosting machines: the classical compute node hosts the NumGenerator, GetGCD, and Controller services, while the quantum compute node hosts QuantumModularExponentiation, QuantumInverseQFT, and Factor services. The hosting of the assembled service is presented as a UML deployment diagram as in Figure 8. UML deployment diagram show two deployment nodes (hosting machines). The classical compute node hosts three services NumGenerator, GetGCD, and Controller, whereas the quantum compute node hosts three services, namely QunatumModularExponentiation, QunatumInverseQFT, and Factor.
    Tool support and architectural language: Figure 8 b) shows the user interface of QADL, a prototype that we developed to model and assemble, i.e., represent and formally describe quantum software components that can provide computational services. Drawing inspiration from classical architectural description languages (ADLs) [67], the quantum architecture description language (QADL) incorporates (a) a graphical environment for modeling software components, (ii) automated parsing for syntactic verification of the components, and (iii) execution of compoent services through integration with IBM Qiskit. In Figure 8 b), left-side, the QADL script specifies the architectural description, declaring QuBits and applying the QUGates (Hadamard, Pauli-X, and CNOT), followed by measurement. This description corresponds to the core quantum subroutine of Shor’s algorithm - superposition (via Hadamard), controlled operations for modular exponentiation (via CNOT), and state inversion (via Pauli-X) - leading to the measurement for computing the periodicity.
@@
    QADL script for architectural specification of Shor’s algorithm via a quantum circuit diagram, as in Appendix A - Listing 2.
@@
    In Figure 8 b), towards the right, QADL represents qunatum circuit diagram to model the algorithm. The description and representation with QADL enables Shor’s algorithm to be modeled, visualised, and executed as a service abstraction that supports design at higher-level of software abstraction rather than low-level implementations. Low-level details, such as circuit creation and component execution within QADL is enabled via IBM Qiskit, an open-source quantum software development framework created by IBM [68]. The Qiskit provides the tools to design, simulate, and run quantum programs on both simulators and real quantum hardware, its primary role is to provide a modular and descriptive framework for describing quantum software architectures.
7 Emerging Trends of QCaaS (RQ-3)
@@
We now answer RQ-3 that aims to report the emerging trends, as dimensions of potential future research, on engineering software services for quantum computing. To identify these trends, we specifically reviewed the details about the objectives, evaluation/demonstrations, as well as limitations and future research details in each reviewed study. This helped us to identify a number of trends that can be generally classified into five (5) main areas, detailed in this section. The identified trends presented in this section do not represent a comprehensive list, instead, it highlights some possible future research based on the data from existing literature. To maintain a consistent presentation of results, we have visualised the emerging trends identified in Figure 9 according to the service lifecycle activities [12] [28] as in Table 2. For example, Figure 9 highlights that during the conception of quantum services, an empirical identification and thematic classification of QSRs is an area of future research. QSRs include quality aspects that complement the functional aspects to ensure the required functionality and desired quality of QCaaS as part of quantum domain engineering activity in QSE.
7.1 Process and Human Support for QCaaS
@@
Process-centred engineering is reflected in the service lifecycle, which structures multiple activities, such as service design, development, and delivery, into a coherent, iterative, and incremental process [24]. Although the reviewed studies organise their findings under the SOA life cycle, most lack an explicitly process-centric approach, thereby underscoring the need for processes tailored to QSE [24]. Emerging paradigms such as quantum DevOps, quantum microservicing, and agile methods for quantum services/software represent quantum-specific adaptations of established software engineering processes that align with the unique requirements of QSE [69]. For example, in an agile approach to QSE, quantum domain engineering can capture domain-specific knowledge of quantum systems (both hardware and software) to develop design models that serve as blueprints for implementing quantum software and services. Process-centric approaches can also support tools (for automation) and professional roles (human decision support) to engineer quantum services. We identified the need for human roles as QSE professionals to effectively undertake QCaaS development. Consequently, there is a clear need to define and develop QSE specific professionals expertise such as quantum architects, domain engineers, and service developers, capable of effectively executing Quantum Computing as a Service (QCaaS) development processes.
@@
Human roles play a pivotal part in enriching the process-centric approach by synergising the theoretical knowledge of quantum physics with the practical disciplines of software engineering. This interdisciplinary integration enables effective support for key activities such as quantum domain engineering, software architecting, and simulation management, areas that currently lack sufficient expertise within quantum service-orientation. By integrating human expertise across these process-centered activities, the architecting process can better address the complexity of designing, developing, and managing quantum-enabled services, ultimately bridging the gap between quantum theory and software practice.
7.2 Empiricism for Managing the QSRs
@@
The concept of Quantum Significant Requirements (QSRs) builds upon the well-established notion of Architecturally Significant Requirements (ASRs) in traditional software and service engineering [49]. Within quantum domain engineering, QSRs play a crucial role in eliciting and documenting both functional and non-functional (quality) aspects of quantum software systems. While existing research has largely concentrated on the functional dimensions of quantum services, it has often overlooked critical non-functional aspects such as QuBit utilisation, energy efficiency, quantum vendor lock-in, QPU elasticity, and quantum error mitigation. These quality attributes, however, are fundamental as they explicitly influence the design, development, and operationalisation of QCaaS solutions. For example, during the quantum domain engineering activity, the hardware aspects (such as the operations of QuGates) can be systematically mapped to software aspects (such as components and connectors). This mapping enables the splitting of computational tasks between classical and quantum computers by applying the classic–quantum split pattern, thereby ensuring an efficient allocation of processing responsibilities across the hybrid computing environment.
@@
Empirical analysis of available data available via the social coding platforms (e.g., GitHub) grounds theoretical QSRs identified via literary analysis, like Qubit utilisation and error mitigation, in a practical and developer-centric context. By examining real-world projects, such as those on GitHub, research can validate or refine the identified requirements with actual developer priorities. This evidence-based approach transforms QSRs from abstract concepts into a practical taxonomy reflective of real quantum software development needs [26]. Furthermore, empirical identification and classification of the QSRs can lead to the creation of a comprehensive QSR taxonomy. By combining the findings of the literature review with data mined from open-source repositories can capture both the existing and emergent, implicit requirements from developer artifacts like issue trackers of qunatum software systems. Classifying these findings allows for the development of a taxonomical classification of QSRs and identifies critical trade-offs, such as between quantum circuit depth and quantum error mitigation [24].
@@
Quantum significant requirements are rooted in the concept of Architecturally Significant Requirements (ASRs), which aim to identify, classify, and document the functional and quality aspects of quantum software services. QSRs for quantum services extend this notion by complementing functional requirements with quality attributes, ensuring that both the required functionality and the desired quality of service are achieved. Furthermore, empirical mining of open-source and social coding platforms can help validate these theoretically derived QSRs, transforming them into a practical, evidence-based taxonomy that reflects the real-world development challenges encountered in quantum software engineering.
@@ 7.3 Model-driven Quantum Software Servicing
Model-Driven Service Engineering (MDSE) empowers software engineers and architects to use models as high-level abstractions that hide complex, implementation-specific details while providing human-comprehensible visual notations for designing and implementing software services. Through model transformations, MDSE enables the conversion of design models into implementation artifacts (e.g., source code) and validation models (e.g., test cases). Modeling notations such as the Service-Oriented Architecture Modeling Language (SoaML) and the Quantum UML (Q-UML) specification [53] extend traditional modeling approaches by offering metamodels and UML profiles tailored for service specification and design within service-oriented architectures. In the context of quantum software engineering, MDSE can particularly aid novice developers by helping them map algorithmic flows and source code modules to graphical representations, thereby supporting low-code, model-driven development of quantum services.
MDSE applies model-driven engineering principles to quantum computing by using notations like SoaML and Q-UML. It abstracts complex implementations into design-level models driven by Quantum Significant Requirements (QSRs). MDQSD enables automation through model transformations and supports human decision-making via tool-assisted design, promoting efficient and low-code development of quantum services.
@@
7.4 Discovery of Quantum Service Patterns
@@
Pattern-based software service engineering leverages architectural designs, implementation strategies, and best practices that can be reused to deliver software services [28]. While existing solutions employ patterns such as the quantum-classic split and service wrapping, there is little empirical evidence on discovering new patterns and tactics as reusable knowledge [70]. This lack of empiricism limits the reusability of design and implementation insights in quantum service engineering. One promising approach for pattern discovery is mining software repositories or social coding platforms (e.g., GitHub) to extract reusable design knowledge. Pattern-based solutions can complement human expertise, enabling novice developers to abstract the complexities of quantum source code while modularising solutions through well-established patterns.
@@
Empirical Discovery of Patterns Empirically grounded methods, such as mining repositories or social coding platforms for quantum software, can capture reusable design rationale for quantum service engineering. Pattern languages enable quantum software architects and algorithm designers to rely on proven best practices rather than ad-hoc, one-off solutions.
7.5 Continuous Testing and Delivery of Services